OpenStudy (anonymous):

convert the angle in degrees to radians . round to 2 decimal places 138 degrees??

OpenStudy (anonymous):

Since 180° = pi radians, (pi = 3.142 . . . ) 1° = pi/180 radians Hence, 138° = 138/180 x 3.142 radians = 2.41 radians.